# End-to-End Conformal Calibration for Optimization Under Uncertainty

### v1

@@ Appendix A Appendix: Related Work
Various designs for deep learning regression models that provide uncertainty estimates have been proposed in the literature, including Bayesian neural networks [10, 19], Gaussian process regression and deep kernel learning [36, 44, 27], ensembles of models [25], and quantile regression [37], among other techniques. These methods typically only provide heuristic uncertainty estimates that are not necessarily well-calibrated [28].
Post-hoc methods such as isotonic regression [24] or conformal prediction [38] may be used to calibrate the uncertainty outputs of deep learning models. These calibration methods generally treat the model as a black box and scale predicted uncertainty levels so that they are calibrated on a held-out calibration set. Isotonic regression guarantees calibrated outputs in the limit of infinite data, whereas conformal methods provide probabilistic, finite-sample calibration guarantees when the calibration set is exchangeable with (e.g., drawn i.i.d. from the same distribution as) the test data. These calibration methods are generally not included in the model training procedure because they involve non-differentiable operators, such as sorting. However, recent works have proposed differentiable losses [17, 40] that approximate the conformal prediction procedure during training and thus allow end-to-end training of models to output more calibrated uncertainty. As approximations, these methods lose the marginal coverage guarantees that true conformal methods provide. However, such guarantees can be recovered at test time by replacing the approximations with true conformal prediction.
@@
The optimization community has proposed a number of techniques over the years to improve robust decision-making under uncertainty, including stochastic, risk-sensitive, chance-constrained, distributionally robust, and robust optimization (e.g., [6, 39, 30, 35]). These techniques have been applied to a wide range of applications, including energy systems operation [46, 29, 16, 47, 33, 43, 8, 14] and portfolio optimization [21, 34, 9]. In these works, the robust and stochastic optimization methods enable selecting decisions (grid resource dispatches or portfolio allocations) in a manner that is aware of uncertainty, e.g., so an energy system operator can ensure that sufficient generation is available to meet demand even on a cloudy day without much solar generation. Typically, however, the construction of uncertainty sets, estimated probability distributions over uncertain parameters, or ambiguity sets over distributions takes place offline and is unconnected to the eventual decision-making task. Thus, our proposed end-to-end approach allows for simultaneous calibration of uncertainty sets with optimal decision-making.
Appendix B Appendix: Maximizing over the uncertainty set
@@
B.1 Maximizing over a box constraint
Let $[\underline{y},\overline{y}]\subset\mathbb{R}^{n}$ be a box uncertainty set for $y\in\mathbb{R}^{n}$. Then, for any vector $c\in\mathbb{R}^{n}$, the primal linear program
@@ B.1 Maximizing over a box constraint
Since strong duality always holds for linear programs, the optimal values of the primal and dual problems will be equal so long as one of the problems is feasible, e.g., so long as the box $[\underline{y},\overline{y}]$ is nonempty. We can thus incorporate this dual problem into the outer minimization of (1) to yield the non-robust form (4).
B.2 Maximizing over an ellipsoid
@@
$\min_{\nu\in\mathbb{R}}\ \frac{1}{4\nu}c^{\top}\Sigma c+\mu^{\top}c+\nu q\qquad\text{s.t.}\quad\nu>0.$
By Slater’s condition, strong duality holds by virtue of the assumption that $q>0$ (which implies strict feasibility of the primal problem), and thus the primal and dual problems have the same optimal value. Moreover, since $\Sigma$ is positive definite and $q>0$, this problem has a unique optimal solution at $\nu^{\star}=\frac{1}{2\sqrt{q}}\|\Sigma^{1/2}c\|_{2}$. Substituting $\nu^{\star}$ into the dual problem objective yields a dual optimal value of
$\sqrt{q}\|\Sigma^{1/2}c\|_{2}+\mu^{\top}c,$
and incorporating this optimal value into the outer minimization of (1) yields the non-robust form (5).
B.3 Proof of Theorem 1: Maximizing over the sublevel set of a PICNN
@@
We can incorporate this dual problem (11) into the outer minimization of (1) to yield the non-robust form (7). For a more interpretable form of this dual problem, let $\nu^{(i)}$ denote the portion of the dual vector $\nu$ corresponding to the $i$-th block-row of matrix $A$, indexed from 00. That is, $\nu^{(i)}=\nu_{id+1:(i+1)d}$ for $i=0,\dotsc,2L-1$. Furthermore, let $\mu=\nu_{2Ld+1}$ be the last entry of $\nu$. Written out, the dual problem (11) becomes
@@
B.4 Ensuring feasibility of the PICNN maximization problem
As noted at the end of Section 3.2, it may sometimes be the case that the inner maximization problem of (1) is unbounded or infeasible when $\Omega_{\theta}(x)$ is parametrized by a PICNN, since in general, the sublevel sets of the PICNN might be unbounded, or the $q$ selected by the split conformal procedure detailed in Section 3.1 may be sufficiently small that $\Omega_{\theta}(x)=\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}$ is empty for certain inputs $x$. We can address each of these concerns using separate techniques.
Ensuring compact sublevel sets
@@ B.4 Ensuring feasibility of the PICNN maximization problem
Ensuring $\Omega_{\theta}(x)$ is nonempty
@@
[…]
@@
In theory, $q_{\min}$ varies as a function of $\theta$, and it is possible to differentiate through the optimization problem (14) using the methods from Agrawal et al. [1] since the problem is convex and $s_{\theta}$ is assumed to be differentiable w.r.t. $\theta$ almost everywhere. However, to avoid this added complexity, in practice, we treat $q_{\min}$ as a constant. In other words, on inputs $x$ where we have to increase $q$ to $q_{\min}$, we treat $\frac{\partial q}{\partial\theta}=0$.
Appendix C Appendix: Exact differentiable conformal prediction
In this section, we prove how to exactly differentiate through the conformal prediction procedure, unlike the approximate derivative first introduced in [40].
Theorem 2.
@@ Appendix C Appendix: Exact differentiable conformal prediction
[…]
@@
[…]
@@
Appendix D Appendix: Experiment details
Our experiments were conducted across a variety of machines, including private servers and Amazon AWS EC2 instances, ranging from 12-core to 128-core machines. Our ETO experiments benefited from GPU acceleration across a combination of NVIDIA GeForce GTX 1080 Ti, Titan RTX, T4, and A100 GPUs. Our E2E experiments did not use GPU acceleration, due to the lack of GPU support in the cvxpylayers [1] Python package.
In all experiments, we use a batch size of 256 and the Adam [23] optimizer.
For all ETO baseline models, we performed a hyperparameter grid search over learning rates ($10^{-4.5}$, $10^{-4}$, $10^{-3.5}$, $10^{-3}$, $10^{-2.5}$, $10^{-2}$, $10^{-1.5}$) and L2 weight decay values (0, $10^{-4}$, $10^{-3}$, $10^{-2}$). Models were trained for up to 100 epochs with early stopping if there was no improvement in validation loss for 10 consecutive epochs.
D.1 Uncertainty representation
Box uncertainty
@@ D.1 Uncertainty representation
In the two-stage ETO baseline, we first train $h_{\theta}$ to estimate the $\alpha/2$- and $(1-\alpha/2)$-quantiles, so that $[h_{\theta}^{\text{lo}}(x),h_{\theta}^{\text{hi}}(x)]$ represents the centered $(1-\alpha)$-confidence region. Quantile regression [37] is a common method for generating uncertainty sets for scalar predictions by estimating quantiles of the conditional distribution $\mathcal{P}(y\mid x)$. For scalar true label $y$, quantile regression models are commonly trained to minimize pinball loss (a.k.a. quantile loss) where $\beta$ is the quantile level being estimated:
$\operatorname{pinball}_{\beta}(\hat{y},y)=\begin{cases}\beta\cdot(y-\hat{y}),&\text{if }y>\hat{y}\\ (1-\beta)\cdot(\hat{y}-y),&\text{if }y\leq\hat{y},\end{cases}$
@@
Our end-to-end (E2E) box uncertainty models use the same architecture as above, initialized with weights from the the trained ETO model. We found it helpful to use a weighted combination of the task loss and pinball loss during training of the E2E models to improve training stability. In our experiments, we used a weight of 0.9 on the task loss and 0.1 on the pinball loss. The E2E models used the best L2 weight decay from the ETO models, and the learning rate was tuned across $10^{-2}$, $10^{-3}$, and $10^{-4}$.
Ellipsoidal uncertainty
Our ellipsoidal uncertainty model uses a neural network $h_{\theta}$ with 3 hidden layers of 256 units each and ReLU activations with batch-normalization. The output layer has dimension $n+n(n+1)/2$, where dimensions $1:n$ predict the mean $\mu_{\theta}(x)$ and the remaining output dimensions are used to construct a lower-triangular Cholesky factor $L_{\theta}(x)$ of the covariance matrix $\Sigma_{\theta}(x)=L_{\theta}(x)L_{\theta}(x)^{\top}$. We pass the diagonal entries of $L$ through a softplus function to ensure strict positivity, which then ensures $\Sigma_{\theta}(x)$ is positive definite.
@@
$\operatorname{NLL}(h_{\theta}(x),y)=-\ln\operatorname{\mathcal{N}}(y\mid\mu_{\theta}(x),\Sigma_{\theta}(x)),$
@@
For the battery storage problem, we used 64 units per hidden layer. We did not run into any feasibility issues for the PICNN maximization problem, so we did not restrict $V_{L}$ as described in Section B.4, and we set $\epsilon=0$.
For the portfolio optimization problem, we used 128 units per hidden layer. We did run into feasibility issues for the PICNN maximization problem, which we resolved by setting $V_{L}=\mathbf{0}_{1\times n}$ as described in Section B.4. This change alone was sufficient, and we set $\epsilon=0$.
For the ETO baseline, we follow the approach by [26] to give probabilistic interpretation to a PICNN model $s_{\theta}$ via the energy-based model $\hat{\mathcal{P}}(y\mid x)\propto\exp(-s_{\theta}(x,y))$. We train our ETO PICNN models with an approximation to the true NLL loss $\text{NLL}(s_{\theta}(x,y))=-\ln\hat{\mathcal{P}}(y\mid x)$ based on samples from Metropolis-Adjusted Langevin Algorithm (MALA), a Markov Chain Monte Carlo (MCMC) method. We refer readers to our code on GitHub for the specific hyperparameters and implementation details we used.
@@
Our end-to-end (E2E) PICNN uncertainty models use the same architecture as above, initialized with weights from the the trained ETO model. Unlike for box and ellipsoidal uncertainty which used a weighted combination of task loss and NLL loss, our E2E PICNN uncertainty models are trained only with the task loss. Similar to the ETO PICNN model, we also regularize the E2E PICNN. Here, we add a regularization loss of $w_{\text{q}}\cdot q^{2}$, where $w_{\text{q}}$ is a regularization weight and $q$ is the conformal prediction threshold computed in each minibatch of E2E training. This regularization loss term aims to keep $q$ near 0; without this regularization, we found that $q$ tended to grow dramatically over training epochs with poor task loss. In our experiments, we set $w_{\text{q}}=0.01$. The E2E models used the best L2 weight decay from the ETO models, and the learning rate was tuned across $10^{-3}$ and $10^{-4}$.
D.2 Data
Price forecasting for battery storage

### v2

@@ Appendix B Appendix: Related Work
Various designs for deep learning regression models that provide uncertainty estimates have been proposed in the literature, including Bayesian neural networks (blundell_weight_2015; gal_dropout_2016), Gaussian process regression and deep kernel learning (rasmussen_gaussian_2005; wilson_deep_2016; liu_simple_2020), ensembles of models (lakshminarayanan_simple_2017), and quantile regression (romano_conformalized_2019), among other techniques. These methods typically only provide heuristic uncertainty estimates that are not necessarily well-calibrated (nado_uncertainty_2022).
Post-hoc methods such as isotonic regression (kuleshov_accurate_2018) or conformal prediction (shafer_tutorial_2008) may be used to calibrate the uncertainty outputs of deep learning models. These calibration methods generally treat the model as a black box and scale predicted uncertainty levels so that they are calibrated on a held-out calibration set. Isotonic regression guarantees calibrated outputs in the limit of infinite data, whereas conformal methods provide probabilistic, finite-sample calibration guarantees when the calibration set is exchangeable (e.g., drawn i.i.d. from the same distribution) with test data. These calibration methods are generally not included in the model training procedure because they involve non-differentiable operators, such as sorting. However, recent works have proposed differentiable losses (einbinder_training_2022; stutz_learning_2022) that approximate the conformal prediction procedure during training and thus allow end-to-end training of models to output more calibrated uncertainty. As approximations, these methods lose the marginal coverage guarantees that true conformal methods provide. However, such guarantees can be recovered at test time by replacing the approximations with true conformal prediction.
@@
The optimization community has proposed a number of techniques over the years to improve robust decision-making under uncertainty, including stochastic, risk-sensitive, chance-constrained, distributionally robust, and robust optimization (e.g., ben-tal_robust_2009; shapiro_lectures_2009; nemirovski_convex_2007; rahimian_distributionally_2019). These techniques have been applied to a wide range of applications, including energy systems operation (zheng_stochastic_2015; ndrio_pricing_2021; dvorkin_chance-constrained_2020; zhong_chance_2021; poolla_wasserstein_2021; warrington_robust_2012; bertsimas_adaptive_2013; christianson_dispatch-aware_2022) and portfolio optimization (gregory_robust_2011; quaranta_robust_2008; bertsimas_data-driven_2018). In these works, the robust and stochastic optimization methods enable selecting decisions (grid resource dispatches or portfolio allocations) in a manner that is aware of uncertainty, e.g., so an energy system operator can ensure that sufficient generation is available to meet demand even on a cloudy day without much solar generation. Typically, however, the construction of uncertainty sets, estimated probability distributions over uncertain parameters, or ambiguity sets over distributions takes place offline and is unconnected to the eventual decision-making task. Thus, our proposed end-to-end approach allows for simultaneous calibration of uncertainty sets with optimal decision-making.
Appendix C Appendix: Maximizing over the uncertainty set
@@
C.1 Maximizing over a box constraint
Let $\underline{y},\overline{y}]\subset\mathbb{R}^{n}$ be a box uncertainty set for $y\in\mathbb{R}^{n}$. Then, for any vector $c\in\mathbb{R}^{n}$, the primal linear program
@@ C.1 Maximizing over a box constraint
Since strong duality always holds for linear programs, the optimal values of the primal and dual problems will be equal so long as one of the problems is feasible, e.g., so long as the box $\underline{y},\overline{y}]$ is nonempty. We can thus incorporate this dual problem into the outer minimization of (1) to yield the non-robust form (4).
C.2 Maximizing over an ellipsoid
@@
$\min_{\nu\in\mathbb{R}}\ \frac{1}{4\nu}c^{\top}\Sigma c+\mu^{\top}c+\nu q\qquad\text{s.t.}\quad\nu\geq 0.$
By Slater’s condition, strong duality holds by virtue of the assumption that $q>0$ (which implies strict feasibility of the primal problem), and thus the primal and dual problems have the same optimal value. Moreover, since $\Sigma$ is positive definite and $q>0$, this problem has a unique optimal solution at $\nu^{\star}=\frac{1}{2\sqrt{q}}\|L^{\top}c\|_{2}$, where $L$ is the unique lower-triangular Cholesky factor of $\Sigma$ (i.e., $\Sigma=LL^{\top}$). Substituting $\nu^{\star}$ into the dual problem yields
$\sqrt{q}\left\lVert L^{\top}c\right\rVert_{2}+\mu^{\top}c.$
Plugging this into (1) yields the non-robust form (5).
We write the dual objective in terms of the Cholesky factor $L$ because our predictive models for ellipsoidal uncertainty directly output the entries of $L$ (see Appendix˜E). Note, however, that the dual problem solution can be equivalently written in terms of the square-root of $\Sigma$, because
$\left\lVert L^{\top}c\right\rVert_{2}^{2}=c^{\top}LL^{\top}c=c^{\top}\Sigma c=c^{\top}\Sigma^{1/2}\Sigma^{1/2}c=\left\lVert\Sigma^{1/2}c\right\rVert_{2}^{2}.$
C.3 Proof of Theorem 4.1: Maximizing over the sublevel set of a PICNN
@@
We can incorporate this dual problem (12) into the outer minimization of (1) to yield the non-robust form (8). For a more interpretable form of this dual problem, let $\nu^{(i)}$ denote the portion of the dual vector $\nu$ corresponding to the $i$-th block-row of matrix $A$, indexed from 0. That is, $\nu^{(i)}=\nu_{id+1:(i+1)d}$ for $i=0,\dotsc,2L-1$. Furthermore, let $\mu=\nu_{2Ld+1}$ be the last entry of $\nu$. Written out, the dual problem (12) becomes
@@
C.4 Ensuring feasibility of the PICNN maximization problem
As noted at the end of Section˜4, it may sometimes be the case that the inner maximization problem of (1) is unbounded or infeasible when $\Omega_{\theta}(x)$ is parametrized by a PICNN, since in general, the sublevel sets of the PICNN might be unbounded, or the $q$ selected by the split conformal procedure detailed in Section˜3.2 may be sufficiently small that $\Omega_{\theta}(x)=\{\hat{y}\in\mathbb{R}^{n}\mid s_{\theta}(x,\hat{y})\leq q\}$ is empty for certain inputs $x$. We can address each of these concerns using separate techniques.
Ensuring compact sublevel sets.
@@ C.4 Ensuring feasibility of the PICNN maximization problem
Ensuring $\Omega_{\theta}(x)$ is nonempty.
@@
$q=\max\left(\min_{\hat{y}\in\mathbb{R}^{n}}s_{\theta}(x,\hat{y}),\ \textsc{Quantile}(\{s_{\theta}(x_{i},y_{i})\}_{(x_{i},y_{i})\in D_{\text{cal}}},\ 1-\alpha)\right).$
@@
In theory, $q_{\min}$ varies as a function of $\theta$, and it is possible to differentiate through the optimization problem (15) using the methods from agrawal_differentiable_2019 since the problem is convex and $s_{\theta}$ is assumed to be differentiable w.r.t. $\theta$ almost everywhere. However, to avoid this added complexity, in practice, we treat $q_{\min}$ as a constant. In other words, on inputs $x$ where we have to increase $q$ to $q_{\min}$, we treat $\frac{\partial q}{\partial\theta}=0$.
Appendix D Appendix: Exact differentiable conformal prediction
In this section, we prove how to exactly differentiate through the conformal prediction procedure, unlike the approximate derivative first introduced in stutz_learning_2022.
Theorem D.1.
@@ Appendix D Appendix: Exact differentiable conformal prediction
Let $q=\textsc{Quantile}(\{s_{i}\}_{i=1}^{M},\ 1-\alpha)$ where the Quantile function is as defined in Algorithm˜1. That is, $q=s_{(k)}$, where $k:=\left\lceil(M+1)(1-\alpha)\right\rceil\in\{1,\dotsc,M,M+1\}$. If $s_{(k)}$ is unique, then
@@
Now, suppose $\alpha\geq\frac{1}{M+1}$. The Quantile function returns the $k$-th largest value of $\{s_{i}\}_{i=1}^{M}\cup\{\infty\}$. Since we assume $s_{(k)}$ is unique, we have $\frac{\mathrm{d}q}{\mathrm{d}s_{(i)}}=\mathbf{1}[i=k]$. Finally, we have
@@
Appendix E Appendix: Experiment details
Our experiments were conducted across a variety of machines, including private servers and Amazon AWS EC2 instances, ranging from 12-core to 128-core machines. Our ETO experiments benefited from GPU acceleration across a combination of NVIDIA GeForce GTX 1080 Ti, Titan RTX, T4, and A100 GPUs. Our E2E experiments did not use GPU acceleration, due to the lack of GPU support in the cvxpylayers Python package (agrawal_differentiable_2019).
In all experiments, we use a batch size of 256 and the Adam optimizer (Kingma2015). Models were trained for up to 100 epochs with early stopping if there was no improvement in validation loss for 10 consecutive epochs.
For box and ellipsoid ETO baseline models, we performed a hyperparameter grid search over learning rates ($10^{-4.5}$, $10^{-4}$, $10^{-3.5}$, $10^{-3}$, $10^{-2.5}$, $10^{-2}$, $10^{-1.5}$) and L2 weight decay values (0, $10^{-4}$, $10^{-3}$, $10^{-2}$). For PICNN ETO models we performed a hyperparameter grid search over learning rates ($10^{-4}$, $10^{-3}$, $10^{-2}$) and L2 weight decay values ($10^{-4}$, $10^{-3}$, $10^{-2}$).
E.1 Uncertainty representation
Box uncertainty.
@@ E.1 Uncertainty representation
In the two-stage ETO baseline, we first train $h_{\theta}$ to estimate the $\alpha/2$- and $(1-\alpha/2)$-quantiles, so that $h_{\theta}^{\text{lo}}(x),h_{\theta}^{\text{hi}}(x)]$ represents the centered $(1-\alpha)$-confidence region. Quantile regression is a common method for generating uncertainty sets for scalar predictions by estimating quantiles of the conditional distribution $\mathcal{P}(y\mid x)$ (romano_conformalized_2019). For scalar true label $y$, quantile regression models are commonly trained to minimize pinball loss (a.k.a. quantile loss) where $\beta$ is the quantile level being estimated:
[…]
@@
Our end-to-end (E2E) box uncertainty models use the same architecture as above, initialized with weights from the trained ETO model. We found it helpful to use a weighted combination of the task loss and pinball loss during training of the E2E models to improve training stability. In our experiments, we used a weight of 0.9 on the task loss and 0.1 on the pinball loss. The E2E models used the best L2 weight decay from the ETO models, and the learning rate was tuned across $10^{-2}$, $10^{-3}$, and $10^{-4}$.
Ellipsoidal uncertainty.
Our ellipsoidal uncertainty model uses a neural network $h_{\theta}$ with 3 hidden layers of 256 units each and ReLU activations with batch-normalization. The output layer has dimension $n+n(n+1)/2$, where dimensions $1:n$ predict the mean $\mu_{\theta}(x)$ and the remaining output dimensions are used to construct a lower-triangular Cholesky factor $L_{\theta}(x)$ of the covariance matrix $\Sigma_{\theta}(x)=L_{\theta}(x)L_{\theta}(x)^{\top}$. We pass the diagonal entries of $L_{\theta}(x)$ through a softplus function to ensure strict positivity, which then ensures $\Sigma_{\theta}(x)$ is positive definite.
@@
$\operatorname{NLL}(\theta)=\frac{1}{N}\sum_{(x,y)\in D}-\ln\operatorname{\mathcal{N}}(y\mid\mu_{\theta}(x),\Sigma_{\theta}(x)),$
@@
For the battery storage problem, we used 64 units per hidden layer. We did not run into any feasibility issues for the PICNN maximization problem, so we did not restrict $V_{L}$ as described in Section˜C.4, and we set $\epsilon=0$.
For the portfolio optimization problem, we tried 32, 64, and 128 units per hidden layer, finding that 32 units worked best. We did run into feasibility issues for the PICNN maximization problem, which we resolved by setting $V_{L}=\mathbf{0}_{1\times n}$ as described in Section˜C.4. This change alone was sufficient, and we set $\epsilon=0$.
For the ETO baseline, we take inspiration from the approach by lin_how_2023 to give probabilistic interpretation to a PICNN model $s_{\theta}$ via the energy-based model $\hat{\mathcal{P}}_{\theta}(y\mid x)=\frac{1}{Z_{\theta}(x)}\exp(-s_{\theta}(x,y))$ where $Z_{\theta}(x):=\int_{\tilde{y}\in\mathbb{R}^{n}}\exp(-s_{\theta}(x,\tilde{y}))\mathop{}\!\mathrm{d}\tilde{y}$ is the normalizing constant. We train our ETO PICNN models with an approximation to the true NLL loss based on samples from the Metropolis-Adjusted Langevin Algorithm (MALA), a Markov Chain Monte Carlo (MCMC) method. We refer readers to our code for the specific hyperparameters and implementation details we used.
@@
Our end-to-end (E2E) PICNN uncertainty models use the same architecture as above, initialized with weights from the the trained ETO model. Unlike for box and ellipsoidal uncertainty which used a weighted combination of task loss and NLL loss, our E2E PICNN uncertainty models are trained only with the task loss. Similar to the ETO PICNN model, we also regularize the E2E PICNN. Here, we add a regularization loss of $w_{\text{q}}\cdot q^{2}$, where $w_{\text{q}}$ is a regularization weight and $q$ is the conformal prediction threshold computed in each minibatch of E2E training. This regularization loss term aims to keep $q$ near 0; without this regularization, we found that $q$ tended to grow dramatically over training epochs with poor task loss. In our experiments, we set $w_{\text{q}}=0.01$.
The E2E models used the best L2 weight decay from the ETO models. For the battery storage problem, we tested learning rates of $10^{-3}$ and $10^{-4}$. For the portfolio optimization problem, we used a learning rate of $5\times 10^{-3}$.
E.2 Data
Price forecasting for battery storage.